OpenStudy (anonymous):

simplify each product you could use F.O.I.L (k^2-4k+3) (k-2)

OpenStudy (anonymous):

i got through the simple tasks but then they added more to the equation and i got confused oh and tanks for helping me

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Let's reverse the order: (k-2)(k^2-4k+3) First, distribute the k to each term of the quadratic. k*k^2-4k*k+3*k = k^3-4k^2+3k Now, let's distribute the -2 to the quadratic. k^2*-2-4k*-2+3*-2 = -2k^2+8k-6 We add the two. k^3-4k^2+3k-2k^2+8k-6 k^3-6k^2+11k-6

OpenStudy (anonymous):

(k^2-4k+3) (k-2) k^3-2k^2-4k^2+8k+3k-6 k^3-6k^2+11k-6

OpenStudy (anonymous):

ok so we would get the same thing even if we reverse it?

OpenStudy (anonymous):

Yes you would. (k^2-4k+3) (k-2) that revered to (k-2) (k^2-4k+3) would give you the same answer